OpenStudy (anonymous):

How do I find 3/7*14/15 please need help

OpenStudy (anonymous):

answer 2/5

OpenStudy (anonymous):

3/7*14/15 =1/1*2/5 (15 will be cancel with 3 at 5 and 14 with 7 at 2) =2/5 (answer)